### v1

Abstract
We present an exact static spherisymmetric solution for the Brans-Dicke action sourced by a self-gravitating massless Klein-Gordon helicity–0 field. In contrast to the Maxwell electromagnetic field, a Klein-Gordon field possesses an energy-momentum tensor with non-vanishing trace. Upon a Weyl mapping into the Einstein frame, the transformed Brans-Dicke scalar field takes on the role of a “dilaton” coupled with the Klein-Gordon field. Despite this dilatonic coupling, the field equations of the resulting Einstein-Klein-Gordon-dilaton action are fully soluble when employing the harmonic radial coordinate. The exact solution derived herein can serve as a prototype for future Brans-Dicke gravity studies involving trace-carrying matter fields. Notably, in the limit of $\omega\rightarrow\infty$, the Brans-Dicke scalar field exhibits an anomalous behavior of $\mathcal{O}(1/\sqrt{\omega})$ as opposed to $\mathcal{O}\left(1/\omega\right)$. As a consequence, the solution converges to a spacetime configuration of General Relativity sourced by the original Klein-Gordon field and a free scalar field, the latter of which is the ${\cal O}\left(1/\sqrt{\omega}\right)$ “remnant” of the Brans-Dicke scalar field. Furthermore, we provide a formal mathematical proof substantiating these two conclusions. Although the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly has been previously discovered for Brans-Dicke vacuum and Brans-Dicke-Maxwell electrovacuum, our findings establish its prevalence in Brans-Dicke gravity regardless of the trace of the energy-momentum tensor of the source. Taken together, the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly challenges the conventional belief in the $\mathcal{O}(1/\omega)$ signature commonly associated with Brans-Dicke gravity. In particular, it may have implications in improving the relativistic corrections to Newtonian gravity beyond the weak-field parametrized post-Newtonian formalism.
I Motivation
The late-time cosmic accelerated expansion observed in 1998 has ignited a keen interest in modifying General Relativity (GR) to accommodate the enigmatic “dark energy” component [1]. Simultaneously, early-time cosmic inflation scenarios often necessitate extensions beyond GR, for example, the Starobinsky model ${\cal R}+{\cal R}^{2}$ [2]. A very active avenue of modified gravity involves introducing a scalar degree of freedom alongside the well-established tensor components in GR [4, 3]. Scalar degrees of freedom naturally emerge in various theoretical frameworks, including Kaluza-Klein, string theory, and braneworld scenarios [3].
A broad class of scalar-tensor theories is encapsulated by the Bergmann-Wagoner action [5, 6]
$S_{\text{ST}}=\int d^{4}x\frac{\sqrt{-g}}{16\pi}\left[\phi\mathcal{R}-\frac{\omega(\phi)}{\phi}\left(\nabla\phi\right)^{2}-2V(\phi)\right]+S^{(m)}$ (1)
This class encompasses $f({\cal R})$ theories (when $\omega(\phi)\equiv 0$, upon certain transformations to re-express the $\phi$ field in terms of ${\cal R}$) and the pure Brans-Dicke (BD) theory ($\omega(\phi)\equiv\omega$, where $\omega$ is a constant, and additionally, $V(\phi)\equiv 0$). The interplay of the scalar field and the metric tensor is a non-trivial aspect of these theories.
Of crucial importance is the limit of the scalar-tensor action (1) to GR. A necessary—but not sufficient—condition for a recovery of GR is that $\phi$ becomes a positive-definite constant field, viz. $\phi\equiv\bar{\phi}>0$, in which case $8\pi/\bar{\phi}$ plays the role of a Newtonian gravitational constant and $V(\bar{\phi})/\bar{\phi}$ acts like a “cosmological constant”. A common way to drive the field $\phi$ to a constant value is by allowing $\omega(\phi)$ to tend to infinity, for example, via the Damour-Nordtvedt dynamical attractor mechanism [7]. For the simplest scalar-tensor version—BD gravity—in which case the Damour-Nordtvedt mechanism is not applicable, a constant field $\phi$ would be achieved by sending the BD parameter $\omega$ to infinity [8].
The large–$\omega$ limit is subtle and delicate, at least for the pure BD version, however. Although the variation $\left(\nabla\phi\right)^{2}$ dies off as $\omega$ grows to infinity, their combination $\omega\left(\nabla\phi\right)^{2}$ may remain significant. There is no a priori theoretical reason that the “kinetic” term $\omega\left(\nabla\phi/\phi\right)^{2}$ in action (1) must be negligible compared with the Einstein-Hilbert term ${\cal R}$ upon the infinite $\omega$ limit. From the variational principle standpoint, when the action expressed in (1) is extremized, the two said terms are in competition, and in principle, they can attain comparable magnitudes.
This scenario is precisely what happens in BD gravity, revealing a non-convergence behavior that has been extensively explored and documented since the 1990s. Specifically, various authors have established that the infinite $\omega$ limit fails to transition the BD vacuum to a vacuum of GR [9, 10, 11, 12, 13, 14, 15, 16, 18, 19, 22, 20, 17, 21, 23]. Instead, it approaches a GR solution sourced by a free massless scalar field; (note: for static spherisymmetric setup, the limiting vacuum is the Fisher-Janis-Newman-Winicour (FJNW) solution  [^1] The solution has been re-discovered several times, and is also known as the Fisher-Bergmann-Leipnik-Janis-Newman-Winicour-Buchdahl-Wyman (FBLJNWBW) solution [23].). Moreover, the convergence rate in the limit is determined to be ${\cal O}\left(1/\sqrt{\omega}\right)$ instead of ${\cal O}\left(1/\omega\right)$. That is to say, the BD scalar field behaves as
$\phi\simeq\bar{\phi}\,\biggl{[}1+\frac{\stackrel{{\scriptstyle 1}}{{\varphi}}}{\sqrt{\omega}}+{\cal O}\left(\frac{1}{\omega}\right)\biggr{]}$ (2)
with the subleading field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ being non-constant. We shall aptly call $\stackrel{{\scriptstyle 1}}{{\varphi}}$ a “remnant” BD field in the infinite $\omega$ limit.
These findings are two faces of the same coin. Due to the $\stackrel{{\scriptstyle 1}}{{\varphi}}$ associated with the $1/\sqrt{\omega}$ order, the “kinetic” term remains finite, viz.
$\omega\left(\nabla\phi/\phi\right)^{2}\simeq\bigl{(}\ \stackrel{{\scriptstyle 1}}{{\varphi}}\>\bigr{)}^{2}$ (3)
and contributes to the limiting vacuum, causing deviation from the GR vacuum. The role of the “remnant” field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ is crucial: it acts as an (additional) free massless scalar field sourcing the FJNW solution mentioned above.
This $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly, as shall be termed in our present article, has also been confirmed for electrovacuum, namely, the BD action coupled with electromagnetism, known as the Brans-Dicke-Maxwell theory [14, 15, 16, 19, 18, 22]. As the energy-momentum tensor (EMT) of the Maxwell electromagnetic field is traceless, the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly has been asserted to exist for BD gravity sourced by traceless matter (a condition trivially met in pure vacuum).
The question concerning fields with a non-vanishing EMT trace remains in contention, however [14, 15]  [^2] Note: it has been advocated in [14, 15] that the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly should be confined to traceless matter as this type of matter possesses certain conformal symmetry, while trace-carrying matter lacks such symmetry. This line of reasoning is circumstantial, however, as the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly might stem from a cause independent from the conformal symmetry at play.. Normal matter, that comprises stars, in general possesses non-vanishing EMT trace, such as a non-relativistic perfect fluid with an EMT given by $T_{\mu}^{\nu}=\text{diag}\left(-\rho,P,P,P\right)$ where $P\ll\rho$. Consequently, if the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly is proven to hold for trace-carrying matter, it could have implications in observational astronomy [24]. For instance, it may indicate the need to go beyond the Robertson parameter $\gamma$ derived in the weak-field parametrized post-Newtonian (PPN) approximation, given by $\gamma_{\text{PPN}}=\frac{\omega+1}{\omega+2}$, which lacks the ${\cal O}\left(1/\sqrt{\omega}\right)$ hallmark.
We plan to tackle this problem in a two-pronged approach. Firstly, we shall present a mathematical proof of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in BD gravity coupled with matter fields of non-vanishing EMT trace. This proof builds upon preliminary ideas previously ventured by one of us in [25, 24]. We find that the “remnant” field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ continues to function as a free massless scalar field in the infinite $\omega$ limit. Its contribution to the “kinetic” term, as per Eq. (3), and the $1/\sqrt{\omega}$ factor associated with the “remnant” field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ via Eq. (2) persist independently of the EMT trace of the matter source.
Secondly, we shall reinforce this proof by examining a concrete case: the BD theory sourced by a massless Klein-Gordon helicity-0 field. For the static spherisymmetric setup, this theory proves to be soluble, and the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ signature manifests in the final analytical solution. As a by-product, the exact solution obtained herein can also serve as a prototype for future studies involving the BD action coupled with trace-carrying matter fields.
A massless Klein-Gordon scalar field $\Psi$ is characterized by an EMT given by
$T_{\mu\nu}^{\Psi}=\partial_{\mu}\Psi\partial_{\nu}\Psi-\frac{1}{2}g_{\mu\nu}g^{\lambda\rho}\partial_{\lambda}\Psi\partial_{\rho}\Psi$ (4)
carrying a non-zero trace, $T^{\Psi}=-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi$. In contrast, the Maxwell electromagnetic (EM) field has a traceless EMT
$T_{\mu\nu}^{F}=-F_{\mu\lambda}F_{\nu}^{\ \lambda}+\frac{1}{4}g_{\mu\nu}F_{\lambda\rho}F^{\lambda\rho}\ \ \Longrightarrow\ \ T^{F}=0$ (5)
The Brans-Dicke-Maxwell (electro)vacuum was first obtained by Bronnikov [26] by extending the works of Penney and others [28, 29, 27]. The success of Bronnikov’s work relies on a Weyl mapping that brings the action into the Einstein frame, significantly simplifying the gravitation sector while leaving the Maxwell field unaffected (a technical point to be re-explained in this paper).
The coupling of Brans-Dicke gravity with a massless Klein-Gordon field has received relatively little attention in the existing literature Unlike the Maxwell EM field, the Klein-Gordon field, upon the Weyl mapping trick, becomes coupled with the ‘dilatonic’ degree of freedom of the BD action in the Einstein frame, thereby complicating the problem  [^3] It should be noted that a closely related extension, allowing GR to couple with a Maxwell field and a dilation field, has been investigated by Clément et al [30].. At first, this technical complexity may seem like a steep price to pay. Nevertheless, we shall demonstrate in this paper that this obstacle can be overcome, leading to an exact analytical solution.
Our paper is organized as follows. In Section II we review the established ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly for BD vacuum and Brans-Dicke-Maxwell electrovacuum. Section III introduces the Brans-Dicke-Klein-Gordon action, outlines the difficulties associated with the Jordan frame, and derives the field equations in the Einstein frame. Sections IV and V are dedicated to solving the Einstein-frame field equations in the harmonic radial coordinate. Section VI examines the solution at large $\omega$, while Section VII provides a formal proof of the observed ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior. Section VIII concludes and offers an outlook on the implication of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in the PPN formalism. The Appendix contains alternative presentations of the solution in various coordinate choices.
II A recap of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in Brans-Dicke vacuum
Consider the Brans-Dicke (BD) action coupled with an external matter source denoted by $\Psi$ [31, 32]:
$\mathcal{S}_{\text{BD}}$ $=\int d^{4}x\sqrt{-g}\Bigl{[}\frac{1}{16\pi}\Bigl{(}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\Bigr{)}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +L^{(m)}\bigl{(}g_{\alpha\beta};\Psi\bigr{)}\Bigr{]}$ (6)
To respect Einstein’s Equivalence Principle, the matter sector $L_{m}$ does not include the BD scalar field $\phi$ [31]. The corresponding field equations are
$\mathcal{R}_{\mu\nu}$ $=\frac{\omega}{\phi^{2}}\nabla_{\mu}\phi\nabla_{\nu}\phi+\frac{1}{\phi}\nabla_{\mu}\nabla_{\nu}\phi+\frac{8\pi}{\phi}\Bigl{[}T_{\mu\nu}-\frac{\omega+1}{2\omega+3}g_{\mu\nu}T\Bigr{]}$ (7)
$\square\,\phi$ $=\frac{8\pi}{2\omega+3}T$ (8)
Here, the EMT of the matter sector is given by
$T_{\mu\nu}:=\frac{-2}{\sqrt{-g}}\frac{\delta\,\bigl{(}\sqrt{-g}\,L^{(m)}(g_{\alpha\beta};\Psi)\bigr{)}}{\delta g^{\mu\nu}}$ (9)
A widely held belief, as popularized in Ref. [8], is that, since $\square\,\phi\simeq\mathcal{O}(1/\omega)$ by virtue of Eq. (8), the scalar field should exhibit the following behavior
$\phi=\bar{\phi}+\mathcal{O}\text{{\hbox{\left(\frac{1}{\omega}\right)}}}$ (10)
with $\bar{\phi}$ being a non-zero constant. This could be seen by expressing the BD scalar field and the metric components in Taylor series of $1/\omega$ as
$\phi$ $=\bar{\phi}\,\biggl{(}1+\frac{\stackrel{{\scriptstyle 2}}{{\varphi}}}{\omega}+\frac{\stackrel{{\scriptstyle 4}}{{\varphi}}}{\omega^{2}}+\dots\biggr{)}$ (11)
$g_{\mu\nu}$ $=\ \stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}+\,\frac{\stackrel{{\scriptstyle 2}}{{g}}_{\mu\nu}}{\omega}+\,\frac{\stackrel{{\scriptstyle 4}}{{g}}_{\mu\nu}}{\omega^{2}}+\dots$ (12)
where $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ is the limiting (or “remnant”) metric as $\omega\rightarrow\infty$, and $\bigl{\{}\stackrel{{\scriptstyle 2n}}{{\varphi}},\,\stackrel{{\scriptstyle 2n}}{{g}}_{\mu\nu}\bigr{\}}$ are corrective components at order $1/\omega^{n}$, respectively. As long as $\bar{\phi}>0$, the asymptotic behavior in (11) results in
$\frac{\omega}{\phi^{2}}\nabla_{\mu}\phi\nabla_{\nu}\phi\simeq\mathcal{O}\biggl{(}\frac{1}{\omega}\biggr{)}\ \text{and}\ \ \frac{1}{\phi}\nabla_{\mu}\nabla_{\nu}\phi\simeq\mathcal{O}\biggl{(}\frac{1}{\omega}\biggr{)}$ (13)
Consequently, with the aid of (11) and (12), Eq. (7) to the leading order produces
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{\mu\nu}=\frac{8\pi}{\bar{\phi}}\left[\stackrel{{\scriptstyle 0}}{{T}}_{\mu\nu}-\frac{1}{2}\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}\stackrel{{\scriptstyle 0}}{{T}}\right]$ (14)
in which $\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{\mu\nu}$ is the Ricci tensor calculated from $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ and $\stackrel{{\scriptstyle 0}}{{T}}_{\mu\nu}$ is the EMT evaluated at $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$. By virtue of Eq. (14),the “remnant” metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ thus satisfies the $\stackrel{{\scriptstyle 0}}{{T}}_{\mu\nu}-$sourced Einstein field equation of General Relativity (GR), with the factor $8\pi/\bar{\phi}$ assuming the role of Newton’s constant.
The $\mathcal{O}(1/\omega)$ behavior of the BD scalar field has faced challenges over the past thirty years. One concrete example central to this debate is the Brans Class I exact solution—a static spherisymmetric vacuum for BD gravity [32]. Previous investigations of this solution [14, 15, 16, 9, 10, 11, 12, 13, 25, 24, 22, 19, 18, 21, 20, 17, 23] have revealed that, in the limit of $\omega\rightarrow\infty$, (i) its metric configuration does not converge to a vacuum of General Relativity (GR), and (ii) its BD scalar field exhibits behavior in the form
$\phi=\bar{\phi}+\mathcal{O}\biggl{(}\frac{1}{\sqrt{\omega}}\biggr{)}$ (15)
in contrast to the conventionally believed form (10).
The appearance of a square root in Expression (15) may seem perplexing at first, as the BD action (6) and resulting field equations (7)–(8) do not inherently carry any non-analytical term. However, a $\sqrt{\omega}$ term has been lurking underneath, and it can be exposed when expressing the action in the Einstein frame, as pointed out by Faraoni and Côté [16]. To see this, through the following Weyl mapping
$\tilde{g}_{\mu\nu}$ $:=\phi\,g_{\mu\nu};\ \ \ \ \ \tilde{\phi}:=\sqrt{2\omega+3}\,\ln\frac{\phi}{\bar{\phi}}$ (16)
the BD action (6) transforms into a GR action minimally coupled with a free scalar field $\tilde{\phi}$, viz.
$\mathcal{S}_{\text{BD}}=\frac{1}{16\pi}\int d^{4}x\sqrt{-\tilde{g}}\Bigl{[}\tilde{\mathcal{R}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\tilde{\phi}\partial_{\nu}\tilde{\phi}\Bigr{]}$ (17)
While this form of the BD action expressed in the Einstein frame is “regular” (i.e., not involving any non-analytical term), upon returning to the Jordan frame, the $\sqrt{2\omega+3}$ participates, per
$\phi=\bar{\phi}\,\exp\frac{\tilde{\phi}}{\sqrt{2\omega+3}}\simeq\bar{\phi}+\mathcal{O}\biggl{(}\frac{1}{\sqrt{\omega}}\biggr{)}$ (18)
hence resulting in the $\mathcal{O}(1/\sqrt{\omega})$ for the BD scalar field. Consequently, this also leads to
[…]
running counter to the series expansion (12). Moreover, the “remnant” metric can be deduced to be proportional to $\tilde{g}_{\mu\nu}$, viz.
$\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}=\bar{\phi}^{-1}\,\stackrel{{\scriptstyle 0}}{{\tilde{g}_{\mu\nu}}}$ (20)
Moreover, it is important to note that, by virtue of the action (17), $\tilde{g}_{\mu\nu}$ is not a vacuum of GR. Rather, it is a configuration of GR sourced by a free massless scalar field (i.e., $\tilde{\phi}$). For the static spherisymmetric setup, $\tilde{g}_{\mu\nu}$ is known to be the FJNW spacetime [34, 37, 33, 35, 36] which deviates from the Schwarzschild spacetime.
This exposition underscores an essential point, however: the square root term $\sqrt{2\omega+3}$ is intrinsic to the gravitational sector itself. That is to say, the $\sqrt{2\omega+3}$ term has nothing to do with the matter sector. Therefore, it is natural to anticipate that the $\sqrt{2\omega+3}$ term should leave its footprints in a Brans-Dicke theory in general, regardless of the trace of the EMT of the source. This anticipation constitutes the central theme of our present paper, wherein we aim to rigorously substantiate and validate this expectation.
It should be noted that prior challenges to the $\mathcal{O}(1/\omega)$ behavior in BD gravity have predominantly focused on scenarios involving vacuum or electrovacuum [9, 10, 11, 12, 13, 14, 15, 16, 18, 19, 22, 20]. Authors have been cautious about extending these challenges to situations where the EMT of the matter source possesses a non-vanishing trace. It has also been contended that while the convergence to GR and the $\mathcal{O}(1/\omega)$ behavior may falter for traceless matter, they remain valid for trace-carrying matter [14, 15]. An exception was one of the current authors who suggested the generality of the $O(1/\sqrt{\omega})$ behavior for non-vanishing EMT trace [25, 24].
III Brans-Dicke action coupled with a Klein-Gordon field
The central focus of our paper is the Brans-Dicke-Klein-Gordon (BDKG) action, where the Brans-Dicke sector is minimally coupled with a free massless Klein-Gordon (KG) scalar field $\Psi$, given as  [^4] For the sake of clarity, we shall drop the $1/(16\pi)$ factor that is conventionally associated with the gravitation sector.
$\mathcal{S}$ $\simeq\sqrt{-g}\biggl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\biggr{]}$ (21)
The EMT of the KG field is $T_{\mu\nu}:=\partial_{\mu}\Psi\partial_{\nu}\Psi-\frac{1}{2}g_{\mu\nu}(\nabla\Psi)^{2}$ which has a non-vanishing trace $T=-(\nabla\Psi)^{2}$. The field equations are
$\mathcal{R}_{\mu\nu}$ $=\frac{\omega}{\phi^{2}}\nabla_{\mu}\phi\nabla_{\nu}\phi+\frac{1}{\phi}\nabla_{\mu}\nabla_{\nu}\phi$
$+\frac{1}{\phi}\left[\partial_{\mu}\Psi\partial_{\nu}\Psi-g_{\mu\nu}\frac{(\nabla\Psi)^{2}}{2(2\omega+3)}\right]$ (22)
$\square\,\phi$ $=-\frac{(\nabla\Psi)^{2}}{2\omega+3}$ (23)
$\square\,\Psi$ $=0$ (24)
These equations are invariant through the two transformations: (1) $\Psi\rightarrow-\Psi$ and (2) $\Psi\rightarrow\Psi+\text{constant}$.
III.1 Difficulties in the Jordan frame
This subsection reveals the technical difficulties associated with the Jordan frame, and the need to move to the Einstein frame. To see this, let us consider a static spherisymmetric metric in the harmonic radial coordinate, viz.
$ds^{2}=-e^{2\gamma(u)}dt^{2}+e^{-2\gamma(u)}\left[e^{2\alpha(u)}du^{2}+e^{\alpha(u)}d\Omega^{2}\right]$ (25)
which yields
$\sqrt{-g}$ $=e^{-2\gamma(u)+2\alpha(u)}\sin\theta$ (26)
$\sqrt{-g}\,g^{uu}$ $=\sin\theta$ (27)
The relevant Ricci tensor components are
$\mathcal{R}_{tt}$ $=e^{4\gamma-2\alpha}\gamma^{\prime\prime}$ (28)
$\mathcal{R}_{uu}$ $=-\alpha^{\prime\prime}+\frac{1}{2}\alpha^{\prime 2}+\gamma^{\prime\prime}-2\gamma^{\prime 2}$ (29)
$\mathcal{R}_{\theta\theta}$ $=1+e^{-\alpha}\Bigl{(}\gamma^{\prime\prime}-\frac{1}{2}\alpha^{\prime\prime}\Bigr{)}$ (30)
In the harmonic radial coordinate, Eqs. (23) and (24) are straightforward to handle. First, Eq. (24) gives (without loss of generality)  [^5] Recall that for a scalar field $\phi$, $\square\phi=(\sqrt{-g})^{-1}\partial_{\mu}(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\phi)$.
$\Psi(u)=\Psi_{*}u$ (31)
with $\Psi_{*}$ being an integration constant. Next, from (23) and (31), we have
$\phi^{\prime\prime}(u)$ $=-\frac{(\Psi^{\prime}(u))^{2}}{2\omega+3}=-\frac{\Psi_{*}^{2}}{2\omega+3}$ (32)
$\Longrightarrow\ \ \phi(u)$ $=-\frac{\Psi_{*}^{2}}{2(2\omega+3)}u^{2}+Nu+v_{0}$ (33)
with $N$ and $v_{0}$ being integration constants. This simplicity is misleading, however. There is “a steep price to pay”: the operator $\nabla_{\mu}\nabla_{\nu}\phi$ in Eq. (22) typically involves the cumbersome Christoffel symbols. The nine non-zero Christoffel symbols are
${\displaystyle\begin{array}[]{lll}\Gamma_{tu}^{t}=\gamma^{\prime};\\ \Gamma_{tt}^{u}=\gamma^{\prime}e^{4\gamma-2\alpha};&\Gamma_{uu}^{u}=\alpha^{\prime}-\gamma^{\prime};\\ \Gamma_{\theta\theta}^{u}=(\gamma^{\prime}-\frac{1}{2}\alpha^{\prime})e^{-\alpha};&\Gamma_{\varphi\varphi}^{u}=\Gamma_{\theta\theta}^{u}\sin^{2}\theta;\\ \Gamma_{u\theta}^{\theta}=-\gamma^{\prime}+\frac{1}{2}\alpha^{\prime};&\Gamma_{\varphi\varphi}^{\theta}=-\cos\theta\sin\theta;\\ \Gamma_{u\varphi}^{\varphi}=-\gamma^{\prime}+\frac{1}{2}\alpha^{\prime};&\Gamma_{\theta\varphi}^{\varphi}=\frac{\cos\theta}{\sin\theta}\end{array}}$ (34)
The field equations for the metric components boil down to
$\left\{\begin{array}[]{rl}\mathcal{R}_{tt}&=\frac{1}{\phi}\nabla_{t}\nabla_{t}\phi-\frac{1}{\phi}\frac{g_{tt}(\nabla\Psi)^{2}}{2(2\omega+3)}\\ &=-\frac{1}{\phi}\Gamma_{tt}^{u}\phi^{\prime}-\frac{1}{\phi}\frac{g_{tt}g^{uu}\Psi^{\prime 2}}{2(2\omega+3)}\\ \mathcal{R}_{uu}&=\frac{\omega}{\phi^{2}}\phi^{\prime 2}+\frac{1}{\phi}\nabla_{u}\nabla_{u}\phi+\frac{1}{\phi}\left[\Psi^{\prime 2}-\frac{g_{uu}(\nabla\Psi)^{2}}{2(2\omega+3)}\right]\\ &=\frac{\omega}{\phi^{2}}\phi^{\prime 2}+\frac{1}{\phi}\left(\phi^{\prime\prime}-\Gamma_{uu}^{u}\phi^{\prime}\right)+\frac{1}{\phi}\frac{4\omega+5}{2(2\omega+3)}\Psi^{\prime 2}\\ \mathcal{R}_{\theta\theta}&=\frac{1}{\phi}\nabla_{\theta}\nabla_{\theta}\phi-\frac{1}{\phi}g_{\theta\theta}\frac{(\nabla\Psi)^{2}}{2(2\omega+3)}\\ &=-\frac{1}{\phi}\Gamma_{\theta\theta}^{u}\phi^{\prime}-\frac{1}{\phi}\frac{g_{\theta\theta}g^{uu}\Psi^{\prime 2}}{2(2\omega+3)}\end{array}\right.$ (35)
which subsequently lead to
$\left\{\begin{array}[]{l}\gamma^{\prime\prime}=-\frac{1}{\phi}\gamma^{\prime}\phi^{\prime}+\frac{1}{\phi}\frac{\Psi^{\prime 2}}{2(2\omega+3)}\\ -\alpha^{\prime\prime}+\frac{1}{2}\alpha^{\prime 2}+\gamma^{\prime\prime}-2\gamma^{\prime 2}\\ \ \ \ \ =\frac{\omega}{\phi^{2}}\phi^{\prime 2}+\frac{1}{\phi}\left(\phi^{\prime\prime}-(\alpha^{\prime}-\gamma^{\prime})\phi^{\prime}\right)+\frac{1}{\phi}\frac{4\omega+5}{2(2\omega+3)}\Psi^{\prime 2}\\ \ e^{\alpha}+\gamma^{\prime\prime}-\frac{\alpha^{\prime\prime}}{2}=-\frac{1}{\phi}\left(\gamma^{\prime}-\frac{\alpha^{\prime}}{2}\right)\phi^{\prime}+\frac{1}{\phi}\frac{\Psi^{\prime 2}}{2(2\omega+3)}\end{array}\right.$ (36)
Solving for $\alpha(u)$ and $\gamma(u)$ from these equations proves to be unwieldy, rendering the Jordan frame unsuitable for obtaining a metric in analytical form. Following the approach introduced by Bronnikov, we will first transition into the Einstein frame, a strategy that has demonstrated advantages [26].
III.2 Transitioning from Jordan frame to Einstein frame
For the sake of illustration, instead of (21), we shall consider a BD action minimally coupled with a massless KG field $\Psi$ and a Maxwell EM field
$\sqrt{-g}\Bigl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi-F_{\mu\nu}F^{\mu\nu}\Bigr{]}$ (37)
where the Maxwell EM field tensor is $F_{\mu\nu}:=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. Let us define a new field $\theta$ per
$\phi:=e^{\theta}$ (38)
and make a Weyl rescaling
$\tilde{g}_{\mu\nu}$ $=e^{\theta}g_{\mu\nu},\ \ \ \tilde{g}^{\mu\nu}=e^{-\theta}g^{\mu\nu},\ \ \ \sqrt{-\tilde{g}}=e^{2\theta}\sqrt{-g}$ (39)
the Ricci scalar transforms as [38]
$\mathcal{R}=e^{\theta}\Bigl{[}\tilde{\mathcal{R}}+3\tilde{\square}\,\theta-\frac{3}{2}(\tilde{\nabla}\theta)^{2}\Bigr{]}$ (40)
Then
$\sqrt{-g}\phi\mathcal{R}=\sqrt{-\tilde{g}}\Bigl{[}\tilde{\mathcal{R}}+3\tilde{g}^{\mu\nu}\tilde{\nabla}_{\mu}(\partial_{\nu}\theta)-\frac{3}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta\Bigr{]}$ (41)
The middle term in right hand side of this last equation is
$\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\tilde{\nabla}_{\mu}(\partial_{\nu}\theta)$ $=\sqrt{-\tilde{g}}\tilde{\nabla}_{\mu}\left(\tilde{g}^{\mu\nu}\partial_{\nu}\theta\right)$ (42)
$=\partial_{\mu}\left(\sqrt{-g}\tilde{g}^{\mu\nu}\partial_{\nu}\theta\right)$ (43)
which is a total derivative (by making use of the identity $\nabla_{\mu}A^{\mu}=\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}A^{\mu})$). Therefore, we arrive at
$\sqrt{-g}\phi\mathcal{R}$ $=\sqrt{-\tilde{g}}\left[\tilde{\mathcal{R}}-\frac{3}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta\right]$ (44)
The kinetic part of the BD scalar field
$\sqrt{-g}\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$ $=\sqrt{-\tilde{g}}\omega\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta$ (45)
For the matter sectors:
$\sqrt{-g}g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi$ $=\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ (46)
$\sqrt{-g}g^{\mu\nu}g^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}$ $=\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\tilde{g}^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}$ (47)
The full action (37) becomes
$\sqrt{-\tilde{g}}\Bigl{[}\tilde{\mathcal{R}}-\left(\omega+\frac{3}{2}\right)\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta-e^{-\theta}\tilde{g}^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\,$
$-\,\tilde{g}^{\mu\nu}\tilde{g}^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}\Bigr{]}$ (48)
Crucially, the Maxwell EM sector remains untransformed; it remains decoupled from the BD scalar field in the Einstein frame. In contrast, the KG field now forms a coupling with the BD scalar field, where the latter takes on the role of a “dilaton”, represented by the term $e^{-\theta}$. This introduces a level of complexity for the KG field.
III.3 Field equations of Einstein-Klein-Gordon-dilation (EKGD) action
In the rest of this paper, we shall not concern with the Maxwell EM sector in action (48). Instead, we shall focus on the following Einstein-Klein-Gordon-dilaton (EKGD) action:
$\sqrt{-g}\left[\mathcal{R}-g^{\mu\nu}\left(\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)\right]$ (49)
Note that, for this section, we have temporarily suppressed the tildes in (49) as compared with (48), for the sake of clarity. We shall restore the tildes in the later sections. Let us find the field equations for the action (49).
Upon variation, standard textbook derivation yields
$\delta\left(\sqrt{-g}\mathcal{R}\right)=\sqrt{-g}\left(\frac{1}{2}g^{\mu\nu}\mathcal{R}-\mathcal{R}^{\mu\nu}\right)\delta g_{\mu\nu}+\text{total deriv.}$. (50)
For other terms, denoting the symmetric rank-2 tensor $A_{\mu\nu}:=\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ and $A:=g^{\mu\nu}A_{\mu\nu}$ the trace of $A_{\mu\nu}$, we have
$\delta\left(\sqrt{-g}g^{\mu\nu}A_{\mu\nu}\right)$
$=\sqrt{-g}\biggl{(}\frac{1}{2}g^{\alpha\beta}\delta g_{\alpha\beta}A+\delta g^{\mu\nu}A_{\mu\nu}+g^{\mu\nu}\delta A_{\mu\nu}\biggr{)}$ (51)
$=\sqrt{-g}\biggl{(}\frac{1}{2}g^{\alpha\beta}\delta g_{\alpha\beta}A-g^{\mu\alpha}g^{\nu\beta}\delta g_{\alpha\beta}A_{\mu\nu}+g^{\mu\nu}\delta A_{\mu\nu}\biggr{)}$ (52)
$=\sqrt{-g}\left(\frac{1}{2}g^{\mu\nu}A-A^{\mu\nu}\right)\delta g_{\mu\nu}+\sqrt{-g}g^{\mu\nu}\delta A_{\mu\nu}$ (53)
Regarding the tensor $A_{\mu\nu}$:
$\delta A_{\mu\nu}=(\omega+3/2)\,\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)+\delta\left(e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$ (54)
The first term is
$\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)$
$=\nabla_{\mu}(\delta\theta)\partial_{\nu}\theta+\partial_{\mu}\theta\nabla_{\nu}(\delta\theta)$ (55)
$=\left[\nabla_{\mu}\left(\delta\theta\partial_{\nu}\theta\right)-\nabla_{\mu}(\partial_{\nu}\theta)\delta\theta\right]$
$+\left[\nabla_{\nu}\left(\delta\theta\partial_{\mu}\theta\right)-\nabla_{\nu}(\partial_{\mu}\theta)\delta\theta\right]$ (56)
$=\left[\nabla_{\mu}\left(\delta\theta\partial_{\nu}\theta\right)+\nabla_{\nu}\left(\delta\theta\partial_{\mu}\theta\right)\right]-2(\nabla_{\mu}\nabla_{\nu}\theta)\,\delta\theta$ (57)
This leads to
$\sqrt{-g}g^{\mu\nu}\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)$
$=2\sqrt{-g}\nabla_{\mu}\left(\delta\theta\nabla^{\mu}\theta\right)-2\sqrt{-g}\square\,\theta\delta\theta$ (58)
[…]
and
$\delta\left(e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$ $=e^{-\theta}\delta\left(\partial_{\mu}\Psi\partial_{\nu}\Psi\right)-\delta\theta e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\,.$ (60)
Similarly to Eq. (59)
$\sqrt{-g}g^{\mu\nu}e^{-\theta}\delta\left(\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$
$=2e^{-\theta}\partial{}_{\mu}\left(\sqrt{-g}\delta\Psi\nabla^{\mu}\Psi\right)-2\sqrt{-g}e^{-\theta}\square\,\Psi\,\delta\Psi$ (61)
[…]
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -2\sqrt{-g}e^{-\theta}\square\,\Psi\,\delta\Psi$ (62)
leading to
$\sqrt{-g}g^{\mu\nu}\delta\left(e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$
$=\sqrt{-g}e^{-\theta}\left[2(\partial_{\mu}\theta\nabla^{\mu}\Psi-\square\,\Psi)\delta\Psi-(\nabla\Psi)^{2}\,\delta\theta\right]$ (63)
Combining all terms, the variation of the EKGD action (49) is
$\sqrt{-g}\biggl{[}\left(\frac{1}{2}g^{\mu\nu}\mathcal{R}-\mathcal{R}^{\mu\nu}\right)-\left(\frac{1}{2}g^{\mu\nu}A-A^{\mu\nu}\right)\biggr{]}\,\delta g_{\mu\nu}$
$+\Bigl{[}(2\omega+3)\sqrt{-g}\square\,\theta+\sqrt{-g}e^{-\theta}(\nabla\Psi)^{2}\Bigr{]}\,\delta\theta$ (64)
$+2\sqrt{-g}e^{-\theta}(-\partial_{\mu}\theta\nabla^{\mu}\Psi+\square\,\Psi)\,\delta\Psi$
Imposing the extremality on (64), regarding $\delta g_{\mu\nu}$, we obtain
$\mathcal{R}^{\mu\nu}-\frac{1}{2}g^{\mu\nu}\mathcal{R}$ $=A^{\mu\nu}-\frac{1}{2}g^{\mu\nu}A$ (65)
which, upon taking the trace, $\mathcal{R}=A$, results in $\mathcal{R}_{\mu\nu}=A_{\mu\nu}$, or
$\mathcal{R}_{\mu\nu}=\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ (66)
Regarding $\delta\theta$, Eq. (64) yields
$(2\omega+3)\,\square\,\theta=-e^{-\theta}(\nabla\Psi)^{2}$ (67)
Regarding $\delta\Psi$, Eq. (64) yields
$\square\,\Psi=\nabla^{\mu}\Psi\nabla_{\mu}\theta$ (68)
The last two equations can also be neatly cast as
$(2\omega+3)\,e^{-\theta}\,\square\,\theta$ $=-(e^{-\theta}\nabla^{\mu}\Psi)\,(e^{-\theta}\nabla_{\mu}\Psi)$ (69)
$\nabla^{\mu}\left(e^{-\theta}\nabla_{\mu}\Psi\right)$ $=e^{-\theta}\left(-\nabla^{\mu}\theta\nabla_{\mu}\Psi+\square\Psi\right)=0$ (70)
When a non-constant Klein-Gordon (KG) field $\Psi$ is present, Eq. (67) requires that the ’dilaton’ field $\theta$ must also be non-constant. It is important to emphasize that the field $\Psi$ is “self-gravitating”, meaning that it must be determined jointly with $g_{\mu\nu}$ and $\theta$ through the derived field equations.
IV Solving the BDKG field equations
Hereafter, we shall restore the tildes in the field equations (66), (69) and (70) to emphasize that they are in the Einstein frame, viz.:
~$\tilde{\mathcal{R}}_{\mu\nu}=\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ (71)
$(2\omega+3)\,e^{-\theta}\,\tilde{\square}\,\theta$ $=-(e^{-\theta}\tilde{\nabla}^{\mu}\Psi)\,(e^{-\theta}\tilde{\nabla}_{\mu}\Psi)$ (72)
~$\tilde{\nabla}^{\mu}\left(e^{-\theta}\tilde{\nabla}_{\mu}\Psi\right)$ $=0$ (73)
Assuming stationarity and spherisymmetry, we shall work in the harmonic radial coordinate in the Einstein frame, viz.
$d\tilde{s}^{2}$ $=-e^{2\gamma(u)}dt^{2}+e^{-2\gamma(u)}\left[e^{2\alpha(u)}du^{2}+e^{\alpha(u)}d\Omega^{2}\right]$ (74)
$\sqrt{-\tilde{g}}$ $=e^{-2\gamma(u)+2\alpha(u)}\sin\theta$ (75)
$\sqrt{-\tilde{g}}\tilde{g}^{uu}$ $=\sin\theta$ (76)
IV.0.1 Solving for the scalar fields
We shall employ the following identities
$\tilde{g}^{\mu\nu}\tilde{\Gamma}_{\mu\nu}^{\lambda}$ $=\frac{-1}{\sqrt{-\tilde{g}}}\partial_{\mu}(\sqrt{-\tilde{g}}\,\tilde{g}^{\lambda\mu})$ (77)
~$\tilde{\square}\,\theta$ $=\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}(\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\Psi)\,.$ (78)
Regarding Eq. (73):
~$\tilde{\nabla}^{\mu}(e^{-\theta}\tilde{\nabla}_{\mu}\Psi)$
$=\tilde{g}^{\mu\nu}\tilde{\nabla}_{\nu}(e^{-\theta}\tilde{\nabla}_{\mu}\Psi)$ (79)
$=\tilde{g}^{\mu\nu}\left[\partial_{\nu}(e^{-\theta}\partial_{\mu}\Psi)-\tilde{\Gamma}_{\mu\nu}^{\lambda}e^{-\theta}\partial_{\lambda}\Psi\right]$ (80)
$=\tilde{g}^{\mu\nu}\partial_{\nu}(e^{-\theta}\partial_{\mu}\Psi)-\left(\tilde{g}^{\mu\nu}\tilde{\Gamma}_{\mu\nu}^{\lambda}\right)e^{-\theta}\partial_{\lambda}\Psi$ (81)
$=\tilde{g}^{\mu\nu}\partial_{\nu}(e^{-\theta}\partial_{\mu}\Psi)+\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}(\sqrt{-\tilde{g}}\,\tilde{g}^{\lambda\mu})e^{-\theta}\partial_{\lambda}\Psi$ (82)
in which the identity (77) has been employed. As $\theta$ and $\Psi$ are functions of $u$, the second term imposes $\lambda=u$, leaving
$\partial_{\mu}(\sqrt{-\tilde{g}}\,\tilde{g}^{u\mu})=\partial_{u}(\sqrt{-\tilde{g}}\,\tilde{g}^{uu})=0$ (83)
The left-hand-side of Eq. (73) becomes
~$\tilde{\nabla}^{\mu}(e^{-\theta}\tilde{\nabla}_{\mu}\Psi)$ $=\tilde{g}^{uu}\partial_{u}(e^{-\theta}\partial_{u}\Psi)$ (84)
$=e^{2\gamma-2\alpha}\left(e^{-\theta}\Psi^{\prime}\right)^{\prime}$ (85)
which then yields
$\Psi^{\prime}=\Psi_{*}e^{\theta}=\Psi_{*}\phi$ (86)
with $\Psi_{*}$ being an integration constant, and $\phi$ the BD scalar field in Jordan frame. With the aid of (86), Eq. (72) yields
$(2\omega+3)\,e^{-\theta}\tilde{\square}\,\theta$ $=-\tilde{g}^{uu}\left(e^{-\theta}\Psi^{\prime}\right)^{2}$ (87)
$\Longrightarrow\ (2\omega+3)\,e^{-\theta}e^{2\gamma-2\alpha}\theta^{\prime\prime}$ $=-e^{2\gamma-2\alpha}\Psi_{*}^{2}$ (88)
or
$\theta^{\prime\prime}$ $=-\frac{\Psi_{*}^{2}}{2\omega+3}e^{\theta}$ (89)
Multiplying both sides with $d\theta=\theta^{\prime}du$ and integrating
$(\theta^{\prime})^{2}+\frac{2\Psi_{*}^{2}}{2\omega+3}e^{\theta}=4\kappa\in\mathbb{R}$ (90)
The case of $\Psi_{*}=0$ would correspond to the known vacuum BD solution. However, we find it more convenient to keep $\Psi_{*}\neq 0$ and manage to recover the BD case in the different manner, to be clarified momentarily. Equation (90) has a solution
$\theta(u)=\begin{cases}\ {\displaystyle\ln\frac{2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\cosh^{2}\sqrt{\kappa}(u+u_{0})}}&(\omega>-3/2,\,\kappa\geqslant 0)\\ \ {\displaystyle\ln\frac{-2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\sinh^{2}\sqrt{\kappa}(u+u_{0})}}&(\omega<-3/2,\,\kappa\in\mathbb{R})\end{cases}$ (91)
or
$\phi(u)=\begin{cases}\ {\displaystyle\frac{2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\cosh^{2}\sqrt{\kappa}(u+u_{0})}}&(\omega>-3/2,\,\kappa\geqslant 0)\\ \ {\displaystyle\frac{-2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\sinh^{2}\sqrt{\kappa}(u+u_{0})}}&(\omega<-3/2,\,\kappa\in\mathbb{R})\end{cases}$ (92)
from which, $\phi(u)\propto e^{-2\sqrt{\kappa}\left|u\right|}\rightarrow 0$ as $u\rightarrow\pm\infty$.
IV.0.2 Recasting the scalar fields using the limit of $\Psi_{*}\rightarrow 0$
Figure 1: $\phi(u)$ and $-\Psi(u)$ as functions of $u$, for various values of $\Psi_{*}$ as indicated in legend. In these plots, $\omega=0.3$, $\kappa\equiv 0.5$, $\phi_{0}\equiv 1$.
For $\Psi_{*}=0$, Eq. (90) yields two affine solutions, $\theta(u)=\pm 2\sqrt{\kappa}(u-u_{0})$ or
$\phi(u)=\phi_{0}\,e^{\pm 2\sqrt{\kappa}u}$ (93)
We may not directly send $\Psi_{*}$ to zero in Eq. (92), however. Some delicate handling is needed. At each value of $\Psi_{*}$, we have the liberty to choose $u_{0}$ in Eq. (92) such that the asymptotic behavior of $\phi(u)$ for large $u$ matches Eq. (93). In particular, for $\omega>-3/2$, (note: $\kappa\geqslant 0$) consider
${\displaystyle\frac{2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\cosh^{2}\sqrt{\kappa}(u+u_{0})}={\displaystyle\frac{8(2\omega+3)\,\kappa}{\Psi_{*}^{2}e^{2\sqrt{\kappa}u_{0}}}\frac{e^{-2\sqrt{\kappa}u}}{\left(1+e^{-2\sqrt{\kappa}(u+u_{0})}\right)^{2}}}}$ (94)
We then set
$\frac{8(2\omega+3)\kappa}{\Psi_{*}^{2}e^{2\sqrt{\kappa}u_{0}}}\equiv\phi_{0}\in\mathbb{R}^{+}$ (95)
or
$e^{-2\sqrt{\kappa}u_{0}}=\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}$ (96)
bringing Eq. (94) to
$\frac{2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\cosh^{2}\sqrt{\kappa}(u+u_{0})}=\frac{\phi_{0}\,e^{-2\sqrt{\kappa}u}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}\right)^{2}}$ (97)
Similarly for $\omega<-3/2$ (with $e^{-2\sqrt{\kappa}u_{0}}=\frac{-\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}$ if $\kappa\geqslant 0$)  [^6] Note: the case of $\omega<-3/2$ and $\kappa<0$ does not have a limit to the pure BD.
$\frac{-2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\sinh^{2}\sqrt{\kappa}(u+u_{0})}$ $=\frac{-8(2\omega+3)\,\kappa}{\Psi_{*}^{2}e^{2\sqrt{\kappa}u_{0}}}\frac{e^{-2\sqrt{\kappa}u}}{\left(1-e^{-2\sqrt{\kappa}(u+u_{0})}\right)^{2}}$ (98)
$=\frac{\phi_{0}\,e^{-2\sqrt{\kappa}u}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}\right)^{2}}$ (99)
In summary, we have chosen a different parametrization for $\phi(u)$. We restrict ourselves to the case of $\kappa\geqslant 0$ (but no restriction on $\omega$, except that $\omega\neq-3/2$). We have set
$u_{0}=-\frac{1}{2\sqrt{\kappa}}\ln\frac{\Psi_{*}^{2}\phi_{0}}{8\left|2\omega+3\right|\kappa}$ (100)
and brought Eq. (92) to the following form
$\phi(u)=\frac{\phi_{0}\,e^{-2\sqrt{\kappa}u}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}\right)^{2}}$ (101)
which gracefully recovers the pure BD case when $\Psi_{*}\rightarrow 0$, viz. $\phi(u)=\phi_{0}e^{-2\sqrt{\kappa}u}$.
[…]
$\Psi(u)$ $=\frac{4(2\omega+3)\sqrt{\kappa}}{\Psi_{*}\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}\right)}-\frac{4(2\omega+3)\sqrt{\kappa}}{\Psi_{*}}$ (102)
$=-\frac{\Psi_{*}\phi_{0}}{2\sqrt{\kappa}}\frac{e^{-2\sqrt{\kappa}u}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}}$ (103)
In the limit of $\Psi_{*}\rightarrow 0$, the KG field disappears as expected. Figure Fig. 1 illustrates the functions $\phi(u)$ and $\Psi(u)$ for various values of $\Psi_{*}$.
IV.0.3 Solving for the metric components
A great virtue of the Einstein frame over the Jordan frame is that the field equation (71) for the metric tensor does not involve the term $\nabla_{\mu}\nabla_{\nu}\phi$ which typically involves the cumbersome Christoffel symbols. Moreover, the $tt-$ and $\theta\theta-$components of Eq. (66) are straightforward to solve as they are decoupled from the scalar fields.
Recall that for the metric expressed in the harmonic radial coordinate (74), the Ricci tensor components are
~$\tilde{\mathcal{R}}_{tt}$ $=e^{4\gamma-2\alpha}\gamma^{\prime\prime}$ (104)
~$\tilde{\mathcal{R}}_{uu}$ $=-\alpha^{\prime\prime}+\frac{1}{2}\alpha^{\prime 2}+\gamma^{\prime\prime}-2\gamma^{\prime 2}$ (105)
~$\tilde{\mathcal{R}}_{\theta\theta}$ $=1+e^{-\alpha}\Bigl{(}\gamma^{\prime\prime}-\frac{1}{2}\alpha^{\prime\prime}\Bigr{)}$ (106)
Invoking the $tt-$component of Eq. (71) and using (104) leads to
$\gamma(u)=-a\,(u+u_{1})\,.$ (107)
Invoking the $\theta\theta-$component of Eq. (71) and using (106) and (107) yields (without loss of generality)
$\alpha^{\prime\prime}$ $=2e^{\alpha}$ (108)
$\Longrightarrow\ \ \alpha(u)$ $=\frac{\chi}{\sinh^{2}\sqrt{\chi}\,u}\,.$ (109)
We should note that for $\chi\in\mathbb{R^{-}}$ the $sinh$ function is expressible as a $sin$ function per
$\frac{\sinh\sqrt{\chi}\,u}{\sqrt{\chi}}$ $=\frac{\sin\sqrt{-\chi}\,u}{\sqrt{-\chi}}\,,$ (110)
and for $\chi=0$, the $sinh$ function admits the following limit
$\lim_{\chi\rightarrow 0}\frac{\sinh\sqrt{\chi}\,u}{\sqrt{\chi}}=u\,.$ (111)
Next, with the aid of (107) and (109), Eq. (105) produces
~$\tilde{\mathcal{R}}_{uu}=2\left(\chi-a^{2}\right)$ (112)
On the other hand, the right-hand-side of the $uu-$component of Eq. (66) is
$\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$
$=\left(\omega+3/2\right)\theta^{\prime 2}+e^{-\theta}\Psi^{\prime 2}$ (113)
$=\left(\omega+3/2\right)\left(-\frac{2\Psi_{*}^{2}}{2\omega+3}e^{\theta}+4\kappa\right)+e^{-\theta}(\Psi_{*}e^{\theta})^{2}$ (114)
$=2(2\omega+3)\,\kappa$ (115)
From Eqs. (112) and (115), the $uu-$component of Eq. (66) becomes a “constraint” among the parameters, given by
$\chi=a^{2}+(2\omega+3)\,\kappa\,.$ (116)
Therefore, the metric in the Einstein frame is
$d\tilde{s}^{2}$ $=-e^{-2a(u+u_{1})}dt^{2}$
$+\frac{e^{2a(u+u_{1})}\chi^{2}}{\sinh^{4}\sqrt{\chi}\,u}du^{2}+\frac{e^{2a(u+u_{1})}\chi}{\sinh^{2}\sqrt{\chi}\,u}d\Omega^{2}\,.$ (117)
IV.0.4 Removal of $u_{1}$
The parameter $u_{1}$ in (117) can be removed via rescaling of variables and parameters, given by
$u$ $\rightarrow e^{au_{1}}u;\ \ \ t\rightarrow e^{au_{1}}t$ (118)
$\chi$ $\rightarrow e^{-2au_{1}}\chi;\ \ \ a\rightarrow e^{-au_{1}}a$ (119)
$\kappa$ $\rightarrow e^{-2au_{1}}\kappa;\ \ \ \Psi_{*}\rightarrow e^{-au_{1}}\Psi_{*}$ (120)
Upon the rescaling, the parameter $u_{1}$ in Eq. (117) is removed, whereas Eqs. (101), (103), and the “constraint” (116) are unaffected.
V Exact solution of BDKG action in the Jordan frame
Moving back to the Jordan frame, the final solution is (with the convention of $a>0$)  [^7] We reemphasize that when $\chi<0$ the function $\frac{\sqrt{\chi}}{\sinh\sqrt{\chi}u}$ is numerically equal to $\frac{\sqrt{-\chi}}{\sin\sqrt{-\chi}u}$.
$ds^{2}=\frac{1}{\phi(u)}\Bigl{[}-e^{-2au}dt^{2}+e^{2au}\left(\frac{\chi^{2}\,du^{2}}{\sinh^{4}\sqrt{\chi}\,u}+\frac{\chi\,d\Omega^{2}}{\sinh^{2}\sqrt{\chi}\,u}\right)\Bigr{]}\,,$ (121)
the BD scalar field is
$\phi(u)={\displaystyle\frac{\phi_{0}\,\exp\left(-2\sqrt{\frac{\chi-a^{2}}{2\omega+3}}u\right)}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\exp\left(-2\sqrt{\frac{\chi-a^{2}}{2\omega+3}}u\right)\right)^{2}}}\,,$ (122)
and the KG field is
$\Psi(u)=-\frac{\Psi_{*}\phi_{0}}{2\sqrt{\frac{\chi-a^{2}}{2\omega+3}}}\frac{\exp\left(-2\sqrt{\frac{\chi-a^{2}}{2\omega+3}}u\right)}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\exp\left(-2\sqrt{\frac{\chi-a^{2}}{2\omega+3}}u\right)}\,.$ (123)
The two scalar fields expressed above are obviously related per Eq. (86):
$\frac{d\Psi(u)}{du}=\Psi_{*}\,\phi(u)\,.$ (124)
The integration constants entering this solution are $\{a,\,\chi,\,\phi_{0},\,\Psi_{*}\}$.
Finally, we have verified by direct inspection that the solution given in (121)–(123) satisfies the field equations (22)–(24) of the original BDKG action (21) in the Jordan frame.
In “reciprocal” coordinate
The final solution (121)–(123) can also be expressed in a more familiar coordinate. Define $r:=1/u$. The limit $u\rightarrow 0$ would correspond to the spatial infinity $r\rightarrow\infty$. The metric (121) becomes
$ds^{2}$ $=\frac{1}{\phi(r)}\biggl{[}-e^{-\frac{2a}{r}}dt^{2}+e^{\frac{2a}{r}}\left(\frac{\sqrt{\chi}/r}{\sinh\sqrt{\chi}/r}\right)^{4}dr^{2}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +e^{\frac{2a}{r}}\left(\frac{\sqrt{\chi}/r}{\sinh\sqrt{\chi}/r}\right)^{2}r^{2}d\Omega^{2}\biggr{]}\,,$ (125)
the BD scalar field (122) is
$\phi(r)={\displaystyle\frac{\phi_{0}\,e^{-\frac{2\sqrt{\kappa}}{r}}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}e^{-\frac{2\sqrt{\kappa}}{r}}\right)^{2}}}\,,$ (126)
and the KG field (123) is
$\Psi(r)=-\frac{\Psi_{*}\phi_{0}}{2\sqrt{\kappa}}\frac{e^{-\frac{2\sqrt{\kappa}}{r}}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}e^{-\frac{2\sqrt{\kappa}}{r}}}\,,$ (127)
in which $\kappa=(\chi-a^{2})/(2\omega+3)$. When $\Psi_{*}=0$, the $\Psi$ field disappears, whereas Expressions (125) and (126) reproduce the four classes of the Brans solution (Class I if $\chi>0$, Class II if $\chi<0$, Classes III and IV if $\chi=0$).
VI Behaviors at large $\omega$
The exact solution (125)–(127) stands handy for us to examine the large $\omega$ limit.
First of all, the static spherisymmetric metric for BDKG theory is conformally equivalent to the FJNW spacetime
$ds_{\text{FJNW}}^{2}$ $=-e^{-\frac{2a}{r}}dt^{2}+e^{\frac{2a}{r}}\left(\frac{\sqrt{\chi}/r}{\sinh\sqrt{\chi}/r}\right)^{4}dr^{2}$
$\ \ \ \ \ \ \ \ \ \ \ +e^{\frac{2a}{r}}\left(\frac{\sqrt{\chi}/r}{\sinh\sqrt{\chi}/r}\right)^{2}r^{2}d\Omega^{2}$ (128)
in which $\chi$ and $a$ are two independent parameters. Only when $\chi=a^{2}\in\mathbb{R}^{+}$ does the metric (128) become the Schwarzschild metric.
For a given pair of $\{\chi,\,a\}$, the metric of BDKG gravity and the FJNW metric would share a common causal structure. For a given casual structure, viz. keeping $\chi$ and $a$ fixed, the dependence on $\omega$ is encapsulated in the conformal factor, viz. $1/\phi(r)$.
VI.0.1 The $\omega\rightarrow\infty$ behavior of the scalar fields
The Taylor expansion of Eq. (126) in terms of $1/\sqrt{\omega}$ produces
$\phi(u)$ $=\frac{\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}$
$-\frac{\phi_{0}\sqrt{2(\chi-a^{2})}\left(1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{3}}\frac{1}{\sqrt{\omega\,}r}+\mathcal{O}\left(\frac{1}{\omega}\right)$ (129)
That is to say, apart from the constant term $\bar{\phi}:=\phi_{0}/\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}$, the subleading contribution to the $\phi$ field is in the order of $1/\sqrt{\omega}$.
Similarly, the Taylor expansion of Eq. (127) yields
$\Psi(r)$ $=-\frac{\Psi_{*}\phi_{0}\,\sqrt{\omega}}{2\sqrt{2(\chi-a^{2})}\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)}$
$+\frac{\Psi_{*}\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}\frac{1}{r}+\mathcal{O}\left(\frac{1}{\sqrt{\omega}}\right)$ (130)
Note that the leading term is a constant; it induces no physical effect and can be dropped.
VI.0.2 The $\omega\rightarrow\infty$ “remnant” metric
With $\chi$ and $a$ held fixed, $\kappa$ behaves as $1/\sqrt{|\omega|}$. Upon sending $\omega$ to infinity, the BD scalar field thus becomes a constant field. However, the limiting metric is not necessarily a Schwarzschild metric. In the harmonic radial coordinate
$ds^{2}=\frac{1}{\bar{\phi}}\Bigl{[}-e^{-2au}dt^{2}+e^{2au}\left(\frac{\chi^{2}\,du^{2}}{\sinh^{4}\sqrt{\chi}\,u}+\frac{\chi\,d\Omega^{2}}{\sinh^{2}\sqrt{\chi}\,u}\right)\Bigr{]}\,,$ (131)
whereas $\phi(u)\equiv\bar{\phi}:=\phi_{0}/\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}$ and $\Psi(u)\equiv\text{const}$. Remark: a constant $\Psi$ has no physical impact at all (while a constant $\phi$ is physically present in the field equations). This is the metric part of the FJNW solution of GR coupled with a scalar field in the form
$S_{\text{BD}}=\int d^{4}x\sqrt{-g}\,\phi_{0}\biggl{(}\mathcal{R}-g^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi\biggr{)}$ (132)
a point that we will justify in the next section.
VII Formal proof of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly
In [25, 24] one of the present authors introduced a set of ideas aimed at establishing the ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior for matter with non-vanishing EMT trace. This section builds upon those ideas, working towards a comprehensive proof. Departing from the conventional expansion in terms of $1/\omega$, we initiate our proof with the following expansions:
$\phi$ $=\bar{\phi}\,\Bigl{(}1+\frac{\stackrel{{\scriptstyle 1}}{{\varphi}}}{\sqrt{\omega}}+\frac{\stackrel{{\scriptstyle 2}}{{\varphi}}}{\omega}+\dots\Bigr{)}$ (133)
$g_{\mu\nu}$ $=\ \stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}+\text{\,}\frac{\stackrel{{\scriptstyle 1}}{{g}}_{\mu\nu}}{\sqrt{\omega}}+\text{\,}\frac{\stackrel{{\scriptstyle 2}}{{g}}_{\mu\nu}}{\omega}+\dots$ (134)
$\Psi$ $=\ \stackrel{{\scriptstyle 0}}{{\Psi}}+\,\frac{\stackrel{{\scriptstyle 1}}{{\Psi}}}{\sqrt{\omega}}+\,\frac{\stackrel{{\scriptstyle 2}}{{\Psi}}}{\omega}+\dots$ (135)
which supersede Eqs. (11) and (12). The metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ shall be referred to as the “background” metric. The critical questions at hand are two-fold: (i) whether the “background” metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ corresponds to the GR solution with only $\stackrel{{\scriptstyle 0}}{{\Psi}}$ as the matter field, and (ii) whether $\stackrel{{\scriptstyle 1}}{{\varphi}}$ exhibits variation. (Note: A constant value of $\stackrel{{\scriptstyle 1}}{{\varphi}}$ can be formally removed by redefining $\bar{\phi}$.) A non-constant field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ would be sufficient to establish the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ hallmark.
VII.1 Proof for BDKG theory
Let us insert the expansions (133)–(135) into the original action (21). To the leading order, we obtain
$\mathcal{S}_{\text{BDKG}}=\ \stackrel{{\scriptstyle 0}}{{\mathcal{S}}}+\mathcal{O}\left(\frac{1}{\sqrt{\omega}}\right)$ (136)
in which the leading term reads
[…]
The action $\stackrel{{\scriptstyle 0}}{{\mathcal{S}}}$ is effectively a “remnant” of the BDKG action upon sending $\omega$ to infinity. The tensor $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ represents a “background” metric. It effectively describes GR in presence of $\stackrel{{\scriptstyle 0}}{{\Psi}}$, but augmented by the scalar $\stackrel{{\scriptstyle 1}}{{\varphi}}$, which is remnant from the vanishing part of the BD scalar. The term $\bar{\phi}^{-1}$ assumes the role of a Newtonian gravitational constant.
Upon variation, the “remnant” action (137) produces the field equations
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{\mu\nu}$ $=\partial_{\mu}\stackrel{{\scriptstyle 1}}{{\varphi}}\partial_{\nu}\stackrel{{\scriptstyle 1}}{{\varphi}}+\bar{\phi}^{-1}\partial_{\mu}\stackrel{{\scriptstyle 0}}{{\Psi}}\partial_{\nu}\stackrel{{\scriptstyle 0}}{{\Psi}}$ (138)
$\stackrel{{\scriptstyle 0}}{{\square}}\ \stackrel{{\scriptstyle 1}}{{\varphi}}$ $=0$ (139)
$\stackrel{{\scriptstyle 0}}{{\square}}\ \stackrel{{\scriptstyle 0}}{{\Psi}}$ $=0$ (140)
Here, the superscript 00 indicates terms and operators evaluated with respect to the background metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$. If a solution to these equations exist for a given $\stackrel{{\scriptstyle 0}}{{\Psi}}$, then this would establish two things:
  1. 1.
    The two non-vanishing terms $\partial_{\mu}\stackrel{{\scriptstyle 0}}{{\Psi}}\partial_{\nu}\stackrel{{\scriptstyle 0}}{{\Psi}}$ and $\partial_{\mu}\stackrel{{\scriptstyle 1}}{{\varphi}}\partial_{\nu}\stackrel{{\scriptstyle 1}}{{\varphi}}$ have independent contributions. Whereas the contribution from $\stackrel{{\scriptstyle 0}}{{\Psi}}$ is expected as a result of the matter source $\Psi$, the contribution from $\stackrel{{\scriptstyle 1}}{{\varphi}}$ is emergent as a result of the BD field dynamics. The latter contribution acts as an additional source to the “background” metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ which then must deviate from a GR solution in the presence of only the KG matter source $\Psi$; and
  2. 2.
    The $\mathcal{O}\left(1/\sqrt{\omega}\right)$ signature in the BD scalar field is self-transparent via $\stackrel{{\scriptstyle 1}}{{\varphi}}$, per Eq. (133).
Specifying to static spherisymmetric setup, it is straightforward to show that the field equations above can be interpreted as an extension of FJNW to two independent scalar fields to have  [^8] We are using the coordinate $r$ which is a reciprocal of the harmonic radial coordinate $u$, viz. $r:=1/u$.
$\stackrel{{\scriptstyle 1}}{{\varphi}}$ $\ =\frac{\sigma}{r}$ (141)
$\stackrel{{\scriptstyle 0}}{{\Psi}}$ $\ =\frac{\Sigma}{r}$ (142)
in which $\sigma$ and $\Sigma$ are “scalar charges”, a term coined by Bronnikov [26], and the “background” metric is given by
$\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}dx^{\mu}dx^{\nu}$ $=-e^{-\frac{2a}{r}}dt^{2}$
$+e^{\frac{2a}{r}}\biggl{(}\frac{(\sqrt{\chi}/r)^{4}}{\sinh^{4}\sqrt{\chi}/r}dr^{2}+\frac{(\sqrt{\chi}/r)^{2}}{\sinh^{2}\sqrt{\chi}/r}\,r^{2}d\Omega^{2}\biggr{)}$ (143)
in which $\chi$, $a$, $\sigma$ and $\Sigma$ are related by
$2\chi=2a^{2}+\sigma^{2}+\bar{\phi}^{-1}\Sigma^{2}$ (144)
Due to the presence of $\stackrel{{\scriptstyle 1}}{{\varphi}}$, the total scalar charge squared gets modified, from $\bar{\phi}^{-1}\Sigma^{2}$ to $\sigma^{2}+\bar{\phi}^{-1}\Sigma^{2}$, affecting the FJNW metric. The BD scalar field $\phi$ thus acquires a behavior
$\phi=\bar{\phi}\,\Bigl{(}1+\frac{\sigma}{\sqrt{\omega}\,r}+\dots\Bigr{)}$ (145)
This completes our proof. When $\sigma=0$ we recover the limiting GR with the original massless Klein-Gordon field as its only source.
VII.2 Verification
The result of limiting $\omega\rightarrow\infty$ has been obtained without resorting to the exact solution. It is desirable to confirm the result using the exact solution. From (129) we extract
$\bar{\phi}\equiv\frac{\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}$ (146)
and
$\stackrel{{\scriptstyle 1}}{{\varphi}}\ =-\frac{1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}\frac{\sqrt{2(\chi-a^{2})}}{r}$ (147)
Similarly, from (130)
$\stackrel{{\scriptstyle 0}}{{\Psi}}\ =\frac{\Psi_{*}\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}\frac{1}{r}$ (148)
the function form of which is consistent with (141) and (142). The scalar charges are
$\sigma$ $=-\frac{1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}\sqrt{2(\chi-a^{2})}$ (149)
$\Sigma$ $=\frac{\Psi_{*}\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}$ (150)
Then, from (146), (149) and (150), it is straightforward to verify that
$\sigma^{2}+\bar{\phi}^{-1}\Sigma^{2}$ $=2(\chi-a^{2})$ (151)
in perfect agreement with (144). Hence, the exact solution is compatible with the proof produced in the preceding subsection.
VII.3 Generalizing to a generic matter source
The proof presented in Section VII.1 can be extended beyond the massless KG field, however. Let us consider a BD action coupled with a matter field $\Psi$ in the form
$\mathcal{S}_{\text{generic}}$ $=\int d^{4}x\sqrt{-g}\Bigl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +L^{(m)}\bigl{(}\{g_{\alpha\beta},\phi\};\Psi\bigr{)}\Bigr{]}$ (152)
The matter sector can be of generic nature, scalar, spinor, massive, etc. It may also couple with the BD scalar field and/or couple with the metric tensor in a non-minimal fashion.
Plugging the Taylor expansions of $1/\sqrt{\omega}$ into the action (152)
$\mathcal{S}_{\text{generic}}=\ \stackrel{{\scriptstyle 0}}{{\mathcal{S}}}_{\text{generic}}+\mathcal{O}\left(\frac{1}{\sqrt{\omega}}\right)$ (153)
in which the leading term
[…]
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -L^{(m)}\Bigl{(}\{\stackrel{{\scriptstyle 0}}{{g}}_{\alpha\beta},\bar{\phi}\};\stackrel{{\scriptstyle 0}}{{\Psi}}\Bigr{)}\biggr{]}$ (154)
Varying the action (154) results in the field equations
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{\mu\nu}\,$ $=\partial_{\mu}\stackrel{{\scriptstyle 1}}{{\varphi}}\partial_{\nu}\stackrel{{\scriptstyle 1}}{{\varphi}}+\bar{\phi}^{-1}\left[\,\stackrel{{\scriptstyle 0}}{{T}}_{\mu\nu}-\,\frac{1}{2}\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}\,\stackrel{{\scriptstyle 0}}{{T}}\,\right]$ (155)
$\stackrel{{\scriptstyle 0}}{{\square}}\ \stackrel{{\scriptstyle 1}}{{\varphi}}\,$ $=0$ (156)
and a conservation law for the EMT (not shown here). The EMT of the matter sector is defined as
$\stackrel{{\scriptstyle 0}}{{T}}_{\mu\nu}:=\left.\frac{-2}{\sqrt{-g}}\frac{\delta\,\bigl{(}\sqrt{-g}\,L^{(m)}(\{g_{\alpha\beta},\phi\};\Psi)\bigr{)}}{\delta g^{\mu\nu}}\right|_{\left\{\begin{array}[]{l}g_{\mu\nu}=\,\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}\\ \phi=\bar{\phi}\\ \Psi=\,\stackrel{{\scriptstyle 0}}{{\Psi}}\end{array}\right.}$ (157)
which is nothing but the original EMT evaluated at $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$, $\bar{\phi}$, and $\stackrel{{\scriptstyle 0}}{{\Psi}}$. Note that the equation for $\stackrel{{\scriptstyle 1}}{{\varphi}}$ is independent of the trace of the EMT.  [^9] We should emphasize that the sourceless “harmonic” equation, Eq. (156), holds in general without assuming stationarity or spherical symmetry. Despite lacking a source, it allows for variations in $\stackrel{{\scriptstyle 1}}{{\varphi}}$. For instance, in the context of a Minkowski background where $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}=\text{diag}(-1,1,1,1)$, the equation reduces to the d’Alembert equation, describing a free propagating wave for $\stackrel{{\scriptstyle 1}}{{\varphi}}$.
In the absence of the BD field, $\stackrel{{\scriptstyle 1}}{{\varphi}}\ \equiv 0$. The background metric is that of GR coupled with the matter field $\Psi$. In the presence of the BD field, the limiting configuration deviates from GR (coupled with the matter field $\Psi$) due to the contribution from $\stackrel{{\scriptstyle 1}}{{\varphi}}$. If $\stackrel{{\scriptstyle 1}}{{\varphi}}$ is non-constant then the $\mathcal{O}(1/\sqrt{\omega})$ signature is established.
For the static spherisymmetric setup, this statement can be proved. The harmonic equation for $\stackrel{{\scriptstyle 1}}{{\varphi}}$ in the harmonic radial coordinate yields a solution
$\stackrel{{\scriptstyle 1}}{{\varphi}}\ \,=\,\sigma\,u$ (158)
with $\sigma$ being a scalar charge  [^10] It should be commented that in the harmonic radial coordinate, the spatial infinity corresponds to $u=0$. The divergence of $\stackrel{{\scriptstyle 1}}{{\varphi}}$ in Eq. (158) when $u$ goes to infinity means that the solution is diverging at a finite $r$. Thence the possibility of the non–GR behavior, in this static case, is related to the presence of a singularity at a finite $r$.. The non-constant $\stackrel{{\scriptstyle 1}}{{\varphi}}$ thus adds a constant contribution to the $uu-$equation, per
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{uu}\ =\,\sigma^{2}+\bar{\phi}^{-1}\left[\,\stackrel{{\scriptstyle 0}}{{T}}_{uu}-\,\frac{1}{2}\stackrel{{\scriptstyle 0}}{{g}}_{uu}\,\stackrel{{\scriptstyle 0}}{{T}}\,\right]$ (159)
Note that the $tt-$ and $\theta\theta-$equations are not affected. Therefore, the functional form of the metric (in the radial harmonic gauge) is unaffected. This is because the functional forms of $\gamma(u)$ and $\alpha(u)$ are fully determined by
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{tt}$ $\ =\bar{\phi}^{-1}\left[\,\stackrel{{\scriptstyle 0}}{{T}}_{tt}-\,\frac{1}{2}\stackrel{{\scriptstyle 0}}{{g}}_{tt}\,\stackrel{{\scriptstyle 0}}{{T}}\,\right]$ (160)
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{\theta\theta}$ $\ =\bar{\phi}^{-1}\left[\,\stackrel{{\scriptstyle 0}}{{T}}_{\theta\theta}-\,\frac{1}{2}\stackrel{{\scriptstyle 0}}{{g}}_{\theta\theta}\,\stackrel{{\scriptstyle 0}}{{T}}\,\right]$ (161)
The $uu-$equation acts as a “constraint”. The relation between the parameters in $A(u)$ and $C(u)$ is altered due to the presence of $\sigma^{2}$ in the new “constraint”. The limiting configuration $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ thus deviates from its usual one without the BD field.
This completes our proof. Note: The proof presented herein does not depend on the trace of the matter source $\Psi$. Moreover, the signature of $1/\sqrt{\omega}$ is self evident via the existence of $\stackrel{{\scriptstyle 1}}{{\varphi}}$.
VIII Conclusion and outlook
The infinite $\omega$ limit in BD gravity is a nuanced matter, encapsulated by the anomalous ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior identified in BD vacuum and electrovacuum since the 1990s [9, 10, 11, 12, 13, 14, 15, 16, 18, 19, 22, 20, 17, 21, 23]. Gaining a clear understanding of this phenomenon could bear significant implications for scalar-tensor theories at large.
Textbook treatments typically assume a diminishing behavior of the BD scalar field $\phi$ in an ${\cal O}\left(1/\omega\right)$ fashion [8]. Due to this presumed requirement, in the infinite $\omega$ limit, the “kinetic” term $\omega\left(\nabla\phi/\phi\right)^{2}$ in the BD action would be suppressed compared with the Einstein-Hilbert term ${\cal R}$, resulting in a recovery to GR with the same matter content. However, from the variational principle standpoint, this scenario is not a priori guaranteed. As the BD action is extremized, the “kinetic” term and the Einstein-Hilbert term can be in competition, leading to an alternative scenario where the “kinetic” term retains a finite and comparable contribution. This latter outcome necessarily compels the BD field to behave as
$\phi\simeq\bar{\phi}\,\biggl{[}1+\frac{\stackrel{{\scriptstyle 1}}{{\varphi}}}{\sqrt{\omega}}+{\cal O}\left(\frac{1}{\omega}\right)\biggr{]}$ (162)
with the subleading field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ being non-constant, and the “kinetic” term surviving the infinite $\omega$ limit, per $\omega\left(\nabla\phi/\phi\right)^{2}\simeq\bigl{(}\ \stackrel{{\scriptstyle 1}}{{\varphi}}\>\bigr{)}^{2}$.
The proof presented in Section VII solidifies this behavior, elucidating its origin and demonstrating that it persists irrespective of the EMT trace of the matter source.
The “remnant” field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ emerges as a free massless scalar field, leaving its footprints in the limiting vacuum. In the broader context of Bergmann-Wagoner theory, the possibility of such a “remnant” field emerging and, if so, its role in the Damour-Nordtvedt attractor mechanism would deserve further examination.
In BD gravity, the presence of a “non-analytical” square-root-of-$\omega$ term may initially seem surprising. However, such a term, viz. $\sqrt{2\omega+3}$, arises during the transition of the action from the Jordan frame to the Einstein frame via a Weyl mapping. The $\sqrt{\omega}$ signature is inherent in the gravitational sector and is unrelated to the matter sector. This is the key reason underlying its prevalence regardless of the EMT trace of matter. We explained this important point in Section II
For the generic Bergmann-Wagoner action (1), the shift to the Einstein frame, viz. $\sqrt{-\tilde{g}}\left[\tilde{{\cal R}}-\frac{1}{2}\left(\tilde{\nabla}\varphi\right)^{2}...\right]$, entails a Weyl mapping given by
$\varphi=\varphi_{0}-\frac{1}{2}\int_{\phi_{0}}^{\phi}\frac{d\phi}{\phi}\sqrt{2\omega(\phi)+3}$ (163)
which is exclusive to the gravitational sector. Hence, for a given function $\omega(\phi)$, any properties that are result from the $\sqrt{2\omega(\phi)+3}$ term should be universal, independent of the EMT trace of matter.
We demonstrated the validity of our proof through a specific case—the BD action coupled with a massless Klein Gordon scalar field (the EMT of which has non-vanishing trace). In Sections IV and V, we expanded upon Bronnikov’s approach [26] by utilizing the Einstein frame and the harmonic radial coordinate. Despite the complication of a dilatonic coupling between the Einstein-frame BD field and the Klein-Gordon field, the system of field equations prove to be soluble. We obtained a static spherisymmetric solution for this coupled action, which can be useful for future studies of BD theory sourced by trace-carrying matter. For the purpose at hand, the characteristic signature of $1/\sqrt{\omega}$ is distinctly evident in the final solution.
Possible extensions of our solution to include Maxwell EM field are deferred for future research. Related studies in GR, coupled with the Maxwell field and a dilaton field, have been reported in [30].
Outlook #1—Let us stress that the issue tackled in this paper is not just a question of formal mathematics, but that genuine potentially new physical features are at stake, in both the $T=0$ and $T\neq 0$ cases. Indeed, it results from an $\mathcal{O}\left(1/\sqrt{\omega}\right)$ behavior of the BD scalar field that the large $\omega$ Brans-Dicke vacuum spherical solution (i.e. Brans class I if $\omega>-3/2$) does not converge to the GR vacuum spherical solution (i.e. Schwarzschild), but to the GR spherical solution filled by a massless scalar (i.e. the FJNW solution). Especially in the strongly scalarized case, the FJNW solution exhibits drastically specific features with respect to Schwarzschild. Indeed, for a given mass of the field, circular orbits exist for any areal radius (while the areal radius is lower bounded in Schwarzschild), with a diverging orbital frequency measured by a remote observer (while this frequency is upper bounded in Schwarzschild), despite the Einstein gravitational red-shifting [39]. Since the associated gravitational wave (GW) frequency, in the Extreme Mass Ration Inspiral (EMRI) scheme, is twice the orbital one, the potential impact of the effect in the new GW astronomy is obvious. Still in the strongly scalarized case, another qualitative specific features with respect to Schwarzschild concerns the light bending, which experiments a maxima when the impact parameter decreases. This results in the presence of a caustic surface, which yields an infinite amplification of the apparent luminosity of a pointlike star crossing this caustics [40]. Thence, the case of an orbiting star on a properly inclined orbit results in a sequence of luminosity peaks, an effect which does not exist in the Schwarzschild spacetime. In the $T\neq 0$ case, the age of the Universe is not constrained to a fixed value by the present Hubble parameter $H_{0}$ in BD dust filled flat Robertson-Walker cosmology, even in the $\omega\rightarrow 0$ limit, while it is fixed to $2/\left(3H_{0}\right)$ in GR dust filled flat Robertson-Walker cosmology [24]. We plan to visit this issue in the future.
Outlook #2—Whether and/or how the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly manifests in physical situations warrant a separate investigation. In particular, the Robertson parameter $\gamma$ of BD gravity has been derived via the parametrized post-Newtonian (PPN) approximation to be $\gamma_{PPN}=\frac{\omega+1}{\omega+2}$, which lacks the ${\cal O}\left(1/\sqrt{\omega}\right)$ signature. This should prompt a scrutiny of the validity of the PPN result, including the underlying assumptions in its derivation. Notably, a recent study has challenged the weak-field assumption inherent in the PPN approximation [41]. We aim to delve into this matter in a future report [42].
Acknowledgements.
H.K.N. thank Valerio Faraoni for helpfully pointing out a number of crucial references, and Mustapha Azreg-Aïnou, Tao Zhu, Luis A. Correa-Borbonet for stimulating discussions.
—————$\infty$—————–
Appendix A   Exact solution of BDKG action in isotropic coordinate
For many purposes, it can be worth rewriting the exact solution (121)–(123) in isotropic form. Since one is finally interested in the large $\omega$ limit, let us specify to the $\omega>-3/2$ case (which, besides, is the condition for the BD theory to be Ostrogradsky stable [43]).
This is achieved by the following radial coordinate transformation
$\left\{\begin{array}[]{l}\rho=K\frac{e^{\sqrt{\chi}u}+1}{e^{\sqrt{\chi}u}-1}\text{ \ \ $\Longleftrightarrow$\ \ }e^{\sqrt{\chi}u}=\frac{\rho+K}{\rho-K}\\ K>0\\ \rho\in\left[K,+\infty\right]\end{array}\right.$ (164)
Let us remark that $\rho=\infty$ for $u=0$. The solution (121)(123) then achieves the form, rescaling $dt$ and $ds$ by a $2K/\sqrt{\chi}$ factor
$\left\{\begin{array}[]{ll}ds^{2}&=\frac{1}{\phi(\rho)}\biggl{[}-\Bigl{(}\frac{\rho-K}{\rho+K}\Bigr{)}^{\frac{2a}{\sqrt{\chi}}}dt^{2}\\ &+\Bigl{(}\frac{\rho^{2}-K^{2}}{\rho^{2}}\Bigr{)}^{2}\Bigl{(}\frac{\rho-K}{\rho+K}\Bigr{)}^{-\frac{2a}{\sqrt{\chi}}}\bigl{(}d\rho^{2}+\rho^{2}d\Omega^{2}\bigr{)}\biggr{]}\\ \phi(\rho)&=\phi_{0}\frac{\left(\frac{\rho-K}{\rho+K}\right)^{\frac{2\sigma}{\sqrt{\chi}}}}{\biggl{[}1+\frac{\Psi_{\ast}^{2}\phi_{0}}{8(\chi-a^{2})}\left(\frac{\rho-K}{\rho+K}\right)^{\frac{2\sigma}{\sqrt{\chi}}}\biggr{]}^{2}}\\ \Psi(\rho)&=-\frac{\Psi_{\ast}\phi_{0}}{2\sigma}\frac{\left(\frac{\rho-K}{\rho+K}\right)^{\frac{2\sigma}{\sqrt{\chi}}}}{1+\frac{\Psi_{\ast}^{2}\phi_{0}}{8(\chi-a^{2})}\left(\frac{\rho-K}{\rho+K}\right)^{\frac{2\sigma}{\sqrt{\chi}}}}\end{array}\right.$ (165)
where
$\sigma\equiv\sqrt{\frac{\chi-a^{2}}{2\omega+3}}$ (166)
Since a global constant factor in the metric is not of significance (it just corresponds to a proper time’s change of units), one can also rewrite the metric as
$ds^{2}$ $=\frac{\phi_{1}}{\phi(\rho)}\biggl{[}-\Bigl{(}\frac{\rho-K}{\rho+K}\Bigr{)}^{\frac{2a}{\sqrt{\chi}}}dt^{2}$
$+\Bigl{(}\frac{\rho^{2}-K^{2}}{\rho^{2}}\Bigr{)}^{2}\Bigl{(}\frac{\rho-K}{\rho+K}\Bigr{)}^{-\frac{2a}{\sqrt{\chi}}}\bigl{(}d\rho^{2}+\rho^{2}d\Omega^{2}\bigr{)}\biggr{]}$ (167)
with
$\phi_{1}=\phi\left(\rho=\infty\right)=\phi_{0}\,\left(1+\frac{\Psi_{\ast}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{-2}$ (168)
for the metric to achieve the standard form ($ds^{2}=-dt^{2}+d\rho^{2}+\rho^{2}d\Omega^{2}$) at spatial infinity.
Appendix B   Exact solution of BDKG action in “standard” coordinate
There exists a convenient expression for the case of $\chi>0$. Let us make a coordinate transform $u(r)$ such that $-\left(\phi g_{tt}\right)\times\left(\phi g_{rr}\right)=1$, viz.
$e^{-2au}\times e^{2au}\frac{\chi^{2}}{\sinh^{4}\sqrt{\chi}u}\left(\frac{du}{dr}\right)^{2}$ $=1$ (169)
From here we obtain:
$\frac{\chi}{\sinh^{2}\sqrt{\chi}u}du=\epsilon\,dr$ (170)
with $\epsilon$ being a sign. Choosing $\epsilon=-$, we get
$r-r_{0}=\frac{\sqrt{\chi}}{\tanh\sqrt{\chi}u}$ (171)
$\Rightarrow u$ $=\frac{1}{\sqrt{\chi}}\text{arctanh}\frac{\sqrt{\chi}}{r-r_{0}}=\frac{1}{2\sqrt{\chi}}\ln\frac{r-r_{0}+\sqrt{\chi}}{r-r_{0}-\sqrt{\chi}}$ (172)
Note: since $\tanh(\sqrt{\chi}\,u)\in(-1,1)$, we must restrict $r$ to be in the range $(-\infty,r_{0}-\sqrt{\chi})\cup(r_{0}+\sqrt{\chi},+\infty)$. Then
$e^{-2au}=\left(\frac{r-r_{0}-\sqrt{\chi}}{r-r_{0}+\sqrt{\chi}}\right)^{\frac{a}{\sqrt{\chi}}}=\left(1-\frac{2\sqrt{\chi}}{r-r_{0}+\sqrt{\chi}}\right)^{\frac{a}{\sqrt{\chi}}}$ (173)
and
$\frac{e^{2au}\chi}{\sinh^{2}\sqrt{\chi}\,u}$ $=\left(1-\frac{2\sqrt{\chi}}{r-r_{0}+\sqrt{\chi}}\right)^{-\frac{a}{\sqrt{\chi}}}\left((r-r_{0})^{2}-\chi\right)$ (174)
Define $r^{\prime}:=r-r_{0}+\sqrt{\chi}$, making $r^{\prime}\in(-\infty,0)\cup(2\sqrt{\chi},+\infty)$, the metric
$ds^{2}$ $=\frac{1}{\phi(r)}\biggl{[}-\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\frac{a}{\sqrt{\chi}}}dt^{2}$
$+\left(1-\frac{2\sqrt{\chi}}{r}\right)^{-\frac{a}{\sqrt{\chi}}}\left(dr^{2}+\left(1-\frac{2\sqrt{\chi}}{r}\right)\,r^{2}\,d\Omega^{2}\right)\biggr{]}$ (175)
The scalar field (in which $\kappa=\frac{\chi-a^{2}}{2\omega+3}\geqslant 0$)
$\phi(r)={\displaystyle\frac{\phi_{0}\,\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}\right)^{2}}}$ (176)
The KG field
$\Psi(r)=-\frac{\Psi_{*}\phi_{0}}{2\sqrt{\kappa}}\frac{\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\left(1-\frac{2\sqrt{\chi}}{r}\right)^{\sqrt{\frac{\kappa}{\chi}}}}\,.$ (177)
If $a=\sqrt{\chi}$, then $\kappa=0$, forcing $\Psi(r)=0$ everywhere. At the same time, the metric is reduced to the Schwarzschild metric
$ds^{2}$ $=-\left(1-\frac{2\sqrt{\chi}}{r}\right)dt^{2}+\left(1-\frac{2\sqrt{\chi}}{r}\right)^{-1}dr^{2}+r^{2}\,d\Omega^{2}$ (178)